OpenStudy (calculusxy):

Algebra 1 help needed! MEDAL!!!

OpenStudy (calculusxy):

@AloneS

OpenStudy (philcua):

qwerrt cant help

OpenStudy (qwertty123):

What is your question?

OpenStudy (philcua):

shell view for a few min tho

OpenStudy (qwertty123):

okay to start off you with.. Does it go up or down how many in Temperature? f(x) = mx + b

OpenStudy (calculusxy):

so i started to do that

OpenStudy (calculusxy):

\[\large \frac{3080-2900}{50 - 54} = m\]

OpenStudy (calculusxy):

@Qwertty123

alones (alones):

@Qwertty123 you better help me, i'm stuck;~; My name there<3

OpenStudy (qwertty123):

I got called back! brb

OpenStudy (calculusxy):

@Qwertty123 It's fine. I have to go now. Thanks!